# Dynamical symmetry breaking in Georgi-Glashow chiral-gauge theories

### v1

@@ Appendix B Full effective action and flow equations
Furthermore, this comparison also allows us to learn more details about the structure of gauge-fermion theories. The computation here presented also allows us to understand the transformation of QCD like results to the chiral limit. One may start with a Dirac QCD-like basis Gies and Jaeckel (2006); Braun (2012); Goertz et al. (2024) and decouple one of the chiralities. Note that this is merely useful to understand the structural level as the product theories are anomalous. The first consequence is that the scalar-pseudoscalar and vector+axial channels are not generated at all as required the presence of two chiralities, as the diagrams are all proportional to $N_{L}N_{R}$. As seen here, in the QCD-like scenario the ${\cal O}_{({\rm V-A})}$ and ${\cal T}_{({\rm V-A})_{{\rm adj}}}$ channels (analogues to our ${\cal O}_{1}$ and ${\cal T}_{2}$) do not fulfill the lowest-order structure for triggering a resonance. Nonetheless, as discussed in Goertz et al. (2024), in the many-flavor QCD limit with locking dynamics evidences for a further resonant ${\cal T}_{({\rm V-A})_{{\rm adj}}}$ channels and condensates has been found, mainly led by strong fermion dynamics in the (S-P) channel. Moreover, the remaining vector-axial structure is indeed generated leading to the results here obtained. A last remark is in order, in the presence of one single chirality the charge-parity symmetry is not present and terms that would violate it in vector-like theories can appear. For example, the four-fermion interactions now receive contributions of the form
@@ Appendix C A closer look at particular examples
Having set the stage, we now integrate the coupled system of four-fermion and gauge flows for different values of $N_{c}$ and $N_{\textrm{gen}}$, as shown in Figure 3. As detailed in the main text, the emergence of an IR fixed point in the flow of the four-fermion couplings indicates that these are fully determined by gauge dynamics. Furthermore, setting different UV boundary conditions does not affect the IR behavior: the system quickly approaches a partial fixed point, as observed in related contexts Goertz et al. (2024); Braun et al. (2016). However, if the UV boundary conditions are too large, e.g., $\bar{\lambda}_{i,\,{k\to\infty}}\sim\mathcal{O}(10)$, the theory exits the universality class of perturbatively renormalizable gauge-fermion systems and enters the domain of nonperturbatively non-renormalizable gauge-NJL models. In that case, the IR behavior depends on the UV data, reflecting the effective nature of these theories. To avoid such effects, we take $\bar{\lambda}_{i,\,{k\to\infty}}=0$ and $\alpha_{g,\,{k\to\infty}}\simeq 0$.
@@
Due to the self-interaction terms $\boldsymbol{c}_{{\rm C},iii}$, the flow equations take a parabolic form (see Figures 4 and 5), where the direction of the parabola depends on the sign of the coefficient. Starting from the UV, the flow for small $\alpha_{g}$ always admits an IR attractive fixed point, which continuously connects to $\bar{\lambda}_{i,\,{k\to\infty}}=0$ as $\alpha_{g}\to 0$. This behavior is insensitive to the coefficients $\boldsymbol{c}_{{\rm A}},\boldsymbol{c}_{{\rm B}},\boldsymbol{c}_{{\rm C}}$, provided $\alpha_{g}$ remains sufficiently small.
@@
However, if any of the couplings displays a resonant structure (see Section IV) and $\alpha_{g}$ reaches $\alpha^{\textrm{crit}}_{\textrm{SB}}$, the corresponding IR fixed point disappears via a merger and a singularity appears in the RG flow. As shown in the main text, for $N_{c}=5$, only two channels meet this condition: ${\cal O}_{5}$ and ${\cal O}_{7}$.
@@
More precisely, $\alpha^{\textrm{crit}}_{\textrm{SB}}$ is reached first, and the singularity appears at a slightly lower scale, signaling the onset of dynamical symmetry breaking at the scale $k_{\textrm{dSB}}$, see Goertz et al. (2024); Ihssen et al. (2024). Note that the scale where $\alpha^{\textrm{crit}}_{\textrm{SB}}$ is reached and $k_{\textrm{dSB}}$ are not identical. The gap between them can grow if other competing dynamics, such as confinement, become important, particularly when the system lies far from the conformal window.
@@
Moving across theory space, the evolution of the gauge coupling changes, as fermionic contributions decelerate its IR growth. Eventually, as $N_{\textrm{gen}}$ increases, the gauge coupling flows to an IR fixed point that remains below $\alpha^{\textrm{crit}}_{\textrm{SB}}$, and no dynamical scale is generated. This scenario is shown in the right-most panel of Figure 3, where all four-fermion couplings approach scale-invariant values as $k\to 0$. Conversely, slightly decreasing $N_{\textrm{gen}}$ may place the IR fixed point just above $\alpha^{\textrm{crit}}_{\textrm{SB}}$, leading to walking behavior and eventual condensation at $k_{\textrm{dSB}}$, as illustrated in the central panel of Figure 3 and discussed in Goertz et al. (2024); Miransky and Yamawaki (1997); Miransky (1999); Braun and Gies (2006, 2010); Braun et al. (2011). Furthermore, as $k_{\textrm{dSB}}$ is approached, all of the couplings start rapidly growing, see left and middle plots in Figure 3. This is an artifact of the four-fermion computation. When bosonising, even partially only the dominant channel, it is known that the subleading tensor structures are not driven to a singularity, see e.g. Goertz et al. (2024); Jaeckel and Wetterich (2003); Denz et al. (2020); Ihssen et al. (2024).
@@
To better visualize the dynamical symmetry breaking, we show in Figures 4 and 5 the beta functions of the four-fermion couplings plotted against themselves for the case $N_{c}=5$, $N_{\textrm{gen}}=7$ (central panel of Figure 3). The flows are displayed at different RG times, from UV (blue) to IR (red), terminating at $k=k_{\textrm{dSB}}$. Here we see that the fixed-point merger, which as we discussed, is predominantly driven by the lowest-order gauge corrections and to a lesser extent by higher-order corrections via other channels.
@@
Finally, we comment on the signs of the four-fermion couplings defined in (5). These are dictated by the lowest-order contributions shown in Figures 2 and 9. For $N_{c}=5$, all couplings except $\bar{\lambda}_{5}$ and $\bar{\lambda}_{7}$ have positive signs in our convention. This choice is motivated by bosonization arguments, where positive sign dressings correspond to a positive curvature of the mesonic potential in the chirally symmetric regime. These non-resonant channels reflect precisely this behavior. Conversely, the resonant channels exhibit negative sign dressings, which in a naive bosonization scheme would imply a tachyonic instability. However, this is not a concern here, as bosonization itself can introduce additional negative signs. Moreover, similar behavior is observed in QCD, where negative-sign couplings coexist with well-understood IR physics.
@@ Appendix D Gauge beta functions and infrared fixed points
$\;\;\;\;\;-\sum_{i}\left[-23C_{2}(R_{i})^{3}+\left(\frac{4204}{54}-\frac{352}{18}\zeta_{3}\right)C_{2}(G)C_{2}(R_{i})^{2}-\left(\frac{7073}{486}-\frac{656}{18}\zeta_{3}\right)C_{2}(G)^{2}C_{2}(R_{i})\right.$
$\;\;\;\;\;+\left.\left(\frac{39143}{162}-\frac{136}{6}\zeta_{3}\right)\,C_{2}(G)^{3}\right]T(R_{i})N_{R_{i}}+\sum_{i}N_{R_{i}}\left(\frac{512}{18}-\frac{1664}{6}\zeta_{3}\right)\frac{d_{R_{i}A}}{{N_{c}}^{2}-1}$
$\;\;\;\;\;-\sum_{i,j}\left[-\left(\frac{184}{12}-\frac{64}{4}\zeta_{3}\right)C_{2}(R_{i})C_{2}(R_{j})+\left(\frac{304}{108}+\frac{128}{36}\zeta_{3}\right)C_{2}(R_{i})^{2}-\left(\frac{17152}{972}+\frac{448}{36}\zeta_{3}\right)C_{2}(R_{i})C_{2}(G)\right.$
$\;\;\;\;\;-\left.\left(\frac{7930}{324}+\frac{224}{36}\zeta_{3}\right)C_{2}(G)^{2}\right]T(R_{i})T(R_{j})N_{R_{i}}N_{R_{j}}-\sum_{i,j}N_{R_{i}}N_{R_{j}}\left(\frac{704}{36}-\frac{512}{12}\zeta_{3}\right)\frac{d_{R_{i}R_{j}}}{{N_{c}}^{2}-1}$
@@
Anti-Fundamental $N$ $\frac{N^{2}-1}{2N}$ $\frac{1}{2}$
Adjoint $N^{2}-1$ $N$ $N$
@@
  Anti-Fund   Adjoint Antisymmetric
Anti-Fund $\frac{({N}^{6}-7{N}^{4}+24{N}^{2}-18)}{96{N}^{2}}$ $\frac{{N}({N}^{4}+5{N}^{2}-6)}{48}$ $\frac{({N}-2)({N}^{2}-1)({N}(N-6)^{2}(N+6)+72)}{96{N}^{2}}$
Adjoint $\frac{{N}^{2}({N}^{4}+35{N}^{2}-36)}{24}$ $\frac{({N}-2)({N}^{2}-1){N}((N-6)N+24)}{48}$
  Antisymmetric   $\frac{(N-2)(N-1)(N+1)((N-2)N(N((N-12)N+48)+144)-576)}{96N^{2}}$

### v2

@@ Appendix B Full effective action and flow equations
Furthermore, this comparison also allows us to learn more details about the structure of gauge-fermion theories. The computation here presented also allows us to understand the transformation of QCD like results to the chiral limit. One may start with a Dirac QCD-like basis Gies:2005as ; Braun:2011pp ; Goertz:2024dnz and decouple one of the chiralities. Note that this is merely useful to understand the structural level as the product theories are anomalous. The first consequence is that the scalar-pseudoscalar and vector+axial channels are not generated at all as required the presence of two chiralities, as the diagrams are all proportional to $N_{L}N_{R}$. As seen here, in the QCD-like scenario the ${\cal O}_{({\rm V-A})}$ and ${\cal T}_{({\rm V-A})_{{\rm adj}}}$ channels (analogues to our ${\cal O}_{1}$ and ${\cal T}_{2}$) do not fulfill the lowest-order structure for triggering a resonance. Nonetheless, as discussed in Goertz:2024dnz , in the many-flavor QCD limit with locking dynamics evidences for a further resonant ${\cal T}_{({\rm V-A})_{{\rm adj}}}$ channels and condensates has been found, mainly led by strong fermion dynamics in the (S-P) channel. Moreover, the remaining vector-axial structure is indeed generated leading to the results here obtained. A last remark is in order, in the presence of one single chirality the charge-parity symmetry is not present and terms that would violate it in vector-like theories can appear. For example, the four-fermion interactions now receive contributions of the form
@@ Appendix C A closer look at particular examples
Having set the stage, we now integrate the coupled system of four-fermion and gauge flows for different values of $N_{c}$ and $N_{\textrm{gen}}$, as shown in Figure 5. As detailed in the main text, the emergence of an IR fixed point in the flow of the four-fermion couplings indicates that these are fully determined by gauge dynamics. Furthermore, setting different UV boundary conditions does not affect the IR behavior: the system quickly approaches a partial fixed point, as observed in related contexts Goertz:2024dnz ; Braun:2014ata . However, if the UV boundary conditions are too large, e.g., $\bar{\lambda}_{i,\,{k\to\infty}}\sim\mathcal{O}(10)$, the theory exits the universality class of perturbatively renormalizable gauge-fermion systems and enters the domain of nonperturbatively non-renormalizable gauge-NJL models. In that case, the IR behavior depends on the UV data, reflecting the effective nature of these theories. To avoid such effects, we take $\bar{\lambda}_{i,\,{k\to\infty}}=0$ and $\alpha_{g,\,{k\to\infty}}\simeq 0$.
@@
Due to the self-interaction terms $\bm{c}_{{\rm C},iii}$, the flow equations take a parabolic form (see Figures 6 and 7), where the direction of the parabola depends on the sign of the coefficient. Starting from the UV, the flow for small $\alpha_{g}$ always admits an IR attractive fixed point, which continuously connects to $\bar{\lambda}_{i,\,{k\to\infty}}=0$ as $\alpha_{g}\to 0$. This behavior is insensitive to the coefficients $\bm{c}_{{\rm A}},\bm{c}_{{\rm B}},\bm{c}_{{\rm C}}$, provided $\alpha_{g}$ remains sufficiently small.
@@
However, if any of the couplings displays a resonant structure (see Section 4) and $\alpha_{g}$ reaches $\alpha^{\textrm{crit}}_{\textrm{SB}}$, the corresponding IR fixed point disappears via a merger and a singularity appears in the RG flow. As shown in the main text, for $N_{c}=5$, only two channels meet this condition: ${\cal O}_{5}$ and ${\cal O}_{7}$.
@@
More precisely, $\alpha^{\textrm{crit}}_{\textrm{SB}}$ is reached first, and the singularity appears at a slightly lower scale, signaling the onset of dynamical symmetry breaking at the scale $k_{\textrm{dSB}}$, see Goertz:2024dnz ; Ihssen:2024miv . Note that the scale where $\alpha^{\textrm{crit}}_{\textrm{SB}}$ is reached and $k_{\textrm{dSB}}$ are not identical. The gap between them can grow if other competing dynamics, such as confinement, become important, particularly when the system lies far from the conformal window.
@@
Moving across theory space, the evolution of the gauge coupling changes, as fermionic contributions decelerate its IR growth. Eventually, as $N_{\textrm{gen}}$ increases, the gauge coupling flows to an IR fixed point that remains below $\alpha^{\textrm{crit}}_{\textrm{SB}}$, and no dynamical scale is generated. This scenario is shown in the right-most panel of Figure 5, where all four-fermion couplings approach scale-invariant values as $k\to 0$. Conversely, slightly decreasing $N_{\textrm{gen}}$ may place the IR fixed point just above $\alpha^{\textrm{crit}}_{\textrm{SB}}$, leading to walking behavior and eventual condensation at $k_{\textrm{dSB}}$, as illustrated in the central panel of Figure 5 and discussed in Goertz:2024dnz ; Miransky1997 ; Miransky:1998dh ; Braun:2006jd ; Braun:2009ns ; Braun:2010qs . Furthermore, as $k_{\textrm{dSB}}$ is approached, all of the couplings start rapidly growing, see left and middle plots in Figure 5. This is an artifact of the four-fermion computation. When bosonising, even partially only the dominant channel, it is known that the subleading tensor structures are not driven to a singularity, see e.g. Goertz:2024dnz ; Jaeckel:2002rm ; Denz:2019ogb ; Ihssen:2024miv .
@@
To better visualize the dynamical symmetry breaking, we show in Figures 6 and 7 the beta functions of the four-fermion couplings plotted against themselves for the case $N_{c}=5$, $N_{\textrm{gen}}=7$ (central panel of Figure 5). The flows are displayed at different RG times, from UV (blue) to IR (red), terminating at $k=k_{\textrm{dSB}}$. Here we see that the fixed-point merger, which as we discussed, is predominantly driven by the lowest-order gauge corrections and to a lesser extent by higher-order corrections via other channels.
@@
Finally, we comment on the signs of the four-fermion couplings defined in (5). These are dictated by the lowest-order contributions shown in Figures 4 and 4. For $N_{c}=5$, all couplings except $\bar{\lambda}_{5}$ and $\bar{\lambda}_{7}$ have positive signs in our convention. This choice is motivated by bosonization arguments, where positive sign dressings correspond to a positive curvature of the mesonic potential in the chirally symmetric regime. These non-resonant channels reflect precisely this behavior. Conversely, the resonant channels exhibit negative sign dressings, which in a naive bosonization scheme would imply a tachyonic instability. However, this is not a concern here, as bosonization itself can introduce additional negative signs. Moreover, similar behavior is observed in QCD, where negative-sign couplings coexist with well-understood IR physics.
@@ Appendix D Gauge beta functions and infrared fixed points
$\;\;\;\;\;-\sum_{i}\left[-23C_{2}(R_{i})^{3}+\left(\frac{4204}{54}-\frac{352}{18}\zeta_{3}\right)C_{2}(G)C_{2}(R_{i})^{2}\right$.
$\;\;\;\;\;-\left.\left(\frac{7073}{486}-\frac{656}{18}\zeta_{3}\right)C_{2}(G)^{2}C_{2}(R_{i})+\left(\frac{39143}{162}-\frac{136}{6}\zeta_{3}\right)\,C_{2}(G)^{3}\right]T(R_{i})N_{R_{i}}$
$\;\;\;\;\;+\sum_{i}N_{R_{i}}\left(\frac{512}{18}-\frac{1664}{6}\zeta_{3}\right)\frac{d_{R_{i}A}}{{N_{c}}^{2}-1}$
$\;\;\;\;\;-\sum_{i,j}\left[-\left(\frac{184}{12}-\frac{64}{4}\zeta_{3}\right)C_{2}(R_{i})C_{2}(R_{j})+\left(\frac{304}{108}+\frac{128}{36}\zeta_{3}\right)C_{2}(R_{i})^{2}\right$.
$\;\;\;\;\;-\left.\left(\frac{17152}{972}+\frac{448}{36}\zeta_{3}\right)C_{2}(R_{i})C_{2}(G)\right.-\left.\left(\frac{7930}{324}+\frac{224}{36}\zeta_{3}\right)C_{2}(G)^{2}\right]T(R_{i})T(R_{j})N_{R_{i}}N_{R_{j}}$
$\;\;\;\;\;-\sum_{i,j}N_{R_{i}}N_{R_{j}}\left(\frac{704}{36}-\frac{512}{12}\zeta_{3}\right)\frac{d_{R_{i}R_{j}}}{{N_{c}}^{2}-1}$
@@
Here, the sum runs over all representations of the theory in terms of left-handed Weyl fermions. In our setup we have {Ri}={               ,
@@
[…]
=Ngen(Nc4)
[…]
. Table 2 summarizes the values of the dimension, the quadratic Casimir $C_{2}(R)$ and the Dynkin index $T(R)$ for relevant representations of $\mathrm{SU}(N)$, and Table 3 contains the coefficient $d_{R_{i}R_{j}}$ entering in the beta function.
@@
Anti-Fundamental NN $\frac{N^{2}-1}{2N}$ $\frac{1}{2}$
Adjoint $^{2}-1 NN$ NN
@@
  Anti-Fund   Adjoint Antisymmetric Anti-Fund $\frac{({N}^{6}-7{N}^{4}+24{N}^{2}-18)}{96{N}^{2}}$ $\frac{{N}({N}^{4}+5{N}^{2}-6)}{48}$ $\frac{({N}-2)({N}^{2}-1)({N}(N-6)^{2}(N+6)+72)}{96{N}^{2}}$ Adjoint $\frac{{N}^{2}({N}^{4}+35{N}^{2}-36)}{24}$ $\frac{({N}-2)({N}^{2}-1){N}((N-6)N+24)}{48}$   Antisymmetric $\frac{(N-2)(N-1)(N+1)((N-2)N(N((N-12)N+48)+144)-576)}{96N^{2}}$